# Adversarial Robustness of Deep Learning Models for Inland Water Body Segmentation from SAR Images

### v1

@@ 1 Introduction
Water body segmentation is a critical task in remote sensing and geospatial analysis, which caters to diverse and critical applications such as flood mapping and monitoring [1], glacial lake segmentation [2], sedimentation and erosion studies [3], and environmental impact assessments. These tasks often rely on high-resolution satellite imagery, such as Synthetic Aperture Radar (SAR) and consequently, accurate segmentation models to delineate water bodies from surrounding land features. SAR imaging uses energy reflectance to be available through day-night lighting and all-weather conditions. However, SAR sensors do not capture data that can discriminate between water and water-like surfaces [4]. Inland water body segmentation is far more challenging than coastline detection, when discussing land-water segmentation of SAR images, owing to the complex geometry of hydrological structures, such as river basins with several tributaries and distributaries (Figure 1).
@@
Deep learning (DL) models, namely convolutional neural networks (CNNs), outperforms traditional approaches in SAR image segmentation in several scenarios. Particularly, encoder-decoder architectures in CNNs, such as U-Net [5] have demonstrated state-of-the-art performance in land-water segmentation of SAR images. Despite these advances in DL, water body segmentation models continue to face challenges during the learning process, such as the presence of noise, domain shifts in remote sensing data, and the impact of adversarial perturbations. Several works in this area rely primarily on manual annotation to generate ground truth masks [1, 6, 7]. These are prone to several human errors, especially during manual annotation. These inadvertent errors include incorrect labels due to variations in river width, and mislabeling boundary pixels. These errors tend to propagate to the image segmentation models using DL, which calls for testing the robustness of these architectures. This motivates our investigation into potential labeling errors and their simulation. Our approach in this study is the novel usage of morphological noise, followed by a systematic testing of adversarial learning of segmentation models, specifically the U-Net architecture. Morphological noise closely models the incorrectness in labels owing to the complex geometry of the hydrological structures of a large river basin, e.g., the Amazon River Basin in South America.
Adversarial attacks pose a significant threat to the robustness and reliability of machine learning models. These attacks involve intentionally crafted perturbations to input data that cause models to misclassify or produce incorrect outputs, often with changes that are barely discernable to the human eye. In the context of water body segmentation, adversarial attacks can lead to inaccurate delineation of water boundaries, with potentially severe consequences for downstream applications such as flood risk assessment or glacial melt monitoring [8, 9]. In this work, the primary form of corruption that we experiment with is data poisoning attacks, which involve corruption of the training data passed to the model. We introduce a specific type of noise, which is the morphological label noise, that is generated using morphological operations (erosion and dilation) in a controlled manner to introduce corruptions in the data points, thus simulating boundary errors.
@@ 2 Related Work
The identification of water bodies from Earth Observation (EO) images is an important step in many applications. Conventionally, a thresholding-based approach is used for water bodies identification, but the thresholds are identified experimentally based on a subset of data [13]. This lacks the capability to generalize, and thresholds may be different for different subsets. Tools based on spectral indices for river identification [14], along with others based on a mathematical morphological-based approach have also been proposed [15]. However, determining the parameters of path opening and closing is challenging and can affect the results obtained.
@@
Recently, semantic segmentation-based approaches have been utilized to identify water bodies in EO images. Semantic segmentation aims to assign a class label to each individual pixel in an image. In the past decade, deep learning-based approaches have obtained excellent results in semantic segmentation for various applications. For instance, several deep learning based approaches have been utilized for semantic segmentation of water bodies [1], including the usage of networks such as FCN-8s [16], tiramisu [17], and pix2pix [18], which is a GAN which has been adapted for the segmentation task. In these supervised learning based segmentation approaches, the training images must be manually labelled by a human. This is a time-consuming and error prone task. Unsupervised learning methods for segmentation on Sentinel-1 SAR imagery [19] have also been explored for river identification. Other interesting works include a Dynamic Snake Convolution [20], which can be used to segment tubular structures such as blood vessels and roads and may be adapted for river segmentation. However, these segmentation approaches focus mainly on accurately identifying water bodies. In comparison, the proposed work studies the performance of segmentation models (primarily U-Net) in water body identification under adversarial attacks from EO images. Our work for semantic segmentation is inspired by DeepRivWidth [21], which uses two primary segmentation models, namely U-Net [5] and DeepLabV3+ [22] for segmentation of Sentinel-1 SAR images.
@@
There exist multiple datasets for water body identification, along with various sources in the form of shapefiles for water body mapping [23, 24, 25, 26, 27]. Examples include the Sen1Floods11 dataset [23], which uses raw Sentinel-1 imagery, and classifies flood water and permanent water, for flood mapping at a global scale, and the Hydrosheds Dataset [24] which provides hydrographic data products including catchment boundaries, rivers, lakes etc. Other examples include DIVA-GIS [25], which provides comprehensive shapefiles for river basins that can be used for ground truth generation, and MERIT Hydro [26], which provides high-resolution hydrography maps using elevation data and water body datasets such as G1WBM (Global 1-second Water Body Map), GSWO (Global Surface Water Occurrence), and OpenStreetMap.
@@
Many of the previously discussed works use human annotation to generate the ground truth for all the images. Not only is this a laborious and time-consuming task, but it is also highly error-prone. There have been a number of works aimed at studying the error types and their effects on the model performance. Most errors from a human annotator occur at the physical boundaries of class labels in spatial applications, which may be classified as biased or unbiased errors [28]. Biased errors may be repeated under similar conditions by the annotator due to human limitations in comprehension and discernibility of the classes, while unbiased errors stem from circumstances (such as a slip of a hand), and are less likely to be repeated in relabeling.
@@
Adversarial attacks are generally classified into white- and black-box attacks, based on the level of access granted to attacker [31]. In white-box attacks, the attackers obtain all information about the target model, including its structure, parameters, etc, and they generally directly interact with the model. On the other hand, for black-box adversaries, the attacker only needs access to the output information of the target model to launch an attack. Black-box attacks are generally more realistic attacks, as the adversary may not always have access to the internal workings of the system. Data poisoning attacks can be either black- or white-box, as they involve changes to the training data itself. Also, depending on the nature of the changes made, they can be considered either a white- or a black-box change [32].
@@
There have also been several works that specifically target data poisoning attacks. One such work proposes a poisoning technique similar to ours, where the adversary targets labels in a multi-class setting and is constrained under a label-flipping budget [35] which is an idea we employ in our work as well. There also exist targeted poisoning attacks, which target one specific test instance, rather than the entire training set [36]. Our work targets the training data instances via a controlled corruption mechanism, which uses morphological operations, and hence, is different from previous studies.
@@ 3 Method
Figure 3: Examples of Sentinel-1 SAR images of the Amazon River Basin (left) in our dataset along with their generated binary water masks (right).
@@ 3.1 Dataset Generation
We selected Sentinel-1 SAR data for our case study of the Amazon River Basin from platforms like the Sentinels Scientific Data Hub and Copernicus Open Access Hub. The Amazon River Basin is an area of interest due to the large area coverage of 7,000,000 km2, with a widespread complex river system. These river structures provide sufficient variety in the training and testing data, as shown in Figure 1. Also, the physical structures of the river and its channels are of complex geometry, which provides several regions where human error could occur while annotating. The chosen images cover 25,194.73 km2 of the basin. The geographic locations of the images in our dataset are shown in Figures 4 and 5.
@@
    The first step is Orbit File Application.
@@
Finally, we implemented a widely-adopted approach using the NDWI index [27]. The multispectral images from Sentinel-2 were obtained using Google Earth Engine. For each image, we clip out the region corresponding to the range of latitude and longitude covered each image, and then, we calculate the NDWI index. This index is used to identify the water regions from the land regions, using a suitable threshold, for each image. This resulted in highly accurate water masks. Examples of the SAR images and their corresponding masks are shown in Figure 3. The entire curated dataset (with correct ground truths and adversarial examples) is published at https://figshare.com/articles/dataset/Adversarial_Attack_Erosion_and_Dilation_Dataset_for_Amazon_River_Basin/28784405?file=53625587.
@@ 3.2 Models for Segmentation
We trained a baseline model using a vanilla U-Net [5] architecture, using our dataset with the accurate, i.e., uncorrupted, ground truth labels. Our goal with this initial setup was to establish a baseline model to benchmark the performance of the model on our dataset. The U-Net model has an encoder with four levels, starting with 64 filters and doubling the number at each level, reaching 512 filters. After downsampling, the bottleneck layer has two convolutional layers with 1024 filters and a dropout layer. The decoder then upsamples the feature maps back to the original resolution. The decoder also applies residual convolutional blocks to refine feature extraction. We then apply a simple threshold of 0.5 on the output of U-Net, to segregate land and water pixels, and generate a binary mask.
@@ 3.3 Simulation of Adversarial Attacks in Training
We also introduce randomness in the kernel sizes used for corruption, to simulate the real-world non-uniformity in the corruption. The stochastic process of selecting the kernels is essential when we need the model to not be biased in its learning. The chosen size depends on the percentage of white pixels (the pixels with mask value 1) in the image. For the images that have a white pixel percentage below 0.2, we give higher probability for choosing the kernel of size 7$\times$7 and lower probability for those with 5$\times$5 and 3$\times$3. We give the probability of 0.7 to that kernel and 0.15 to each other kernel. The exact opposite happens for dilation, we give high probability for 3$\times$3 and lower probability values for the other two kernels when the white pixel percentage is high. Finally, we utilize the following levels of corruption, given the image sizes are $512\times 512$: $2\%$, $12\%$, $15\%$, $17\%$, $20\%$, $25\%$, $30\%$}.
With the model trained on correct masks, we now introduce controlled corruptions to the masks by using erosion and dilation, as these operations present examples of potential human errors in precise delineation of land-water boundaries. Thus, these controlled corruptions are morphological label noise, and this process of corruption is effectively a data poisoning procedure.
@@ 3.4 Implementation
The data points are divided into 883 training, 127 validation, and 253 test samples. The U-Net model is trained for 20 epochs, for this experiment and across all subsequent experiments. We used a pre-trained implementation of Deeplabv3 with Resnet-50 backbone, and finetuned it over our dataset, for 50 epochs (with an Early Stopping condition to prevent overfitting). We used Segformer (b-4-sized) pretrained on Imagenet [45], and finetuned for 50 epochs on our dataset. Finally, for Mask2former, we used the large-sized version with Swin Transformer, pretrained on the Cityscapes dataset [46], and finetuned it for a total of 50 epochs on the dataset. We run all the models on a system with an AMD Ryzen 9 5950X 16-core processor, connected to NVIDIA RTX A4000 graphics card. Our code is available at https://github.com/GVCL/IWSeg-SAR-Poison.git.
@@ 4.1 Segmentation Results with the Uncorrupted Dataset
We present the results of state-of-the-art segmentation models (U-Net, DeepLabV3, Segformer, Mask2former) on our dataset (uncorrupted), which help us assess the quality of the dataset. Table 1 shows the performance of the models on the dataset. The performance of each of the models saturates at a per pixel accuracy of $\sim$0.955. Moreover, the IoU for all these models are comparable.
We further investigate into the samples that none of the models are able to segment properly. A few such failure cases are discussed in Section 4.3, where either due to incorrect thresholding, or due to unavailability of a complete Sentinel-2 image, the ground truth itself is incorrect. The incorrectness of the ground truth consequently leads to a higher level of segmentation inaccuracy.
However, one consistent failure observed across the model is their inability to fully delineate narrow tributaries in the image. An example of one such failure case is shown in Figure 6. Overall, we observe that all our selected models face issues in delineating the very narrow tributaries, and are only somewhat correct, with varying levels of success.
@@ 4.2 Adversarial Experimentation
This saturation of corruption is demonstrated in Figures 8 and 8. Figure 8 shows the box plot of the actual number of corrupted pixels, across all training images for different corruption levels, where the corruption is by erosion, and a similar box plot for dilation. We use the Structural Similarity Index (SSIM) score to quantify the similarities between two images. Figure 8 shows the distribution of the SSIM scores between the uncorrupted ground truth and the corresponding corrupted ground truth, computed across all images and for different corruption levels. We observe that the number of pixels corrupted for images undergoing erosion becomes stagnant as we increase the corruption levels, with the median and the lower quartile barely changing beyond $20\%$ corruption. The dilation plots, on the other hand, show the expected trends. The same is true for the SSIM values between the uncorrupted masks and their corresponding corrupted masks. Thus, we conclude that, in our approach, the corruptions may be limited by design, as the final count of corrupt pixels for an image may be lesser than that expected in the targeted level of corruption.
@@ 4.3 Results on Corrupted Dataset and Discussion
Table 2 shows the evaluation metrics of training the U-Net model, across the various levels of corruption. We get a good 95% accuracy in segmentation. Furthermore, the performance is mostly unaffected even after increasing the corruption levels to as high as 20%. This is surprising, as the model is able to tolerate even such high levels of corruption. We observe that the model performance does not deteriorate significantly until 25% corruption. Hence, we investigate more into the training of the model, the qualitative predictions of the model on a particular image, and the actual corruption that the model is facing.
@@
Figure 9 shows the per-epoch accuracy on the training and validation data for each of the corruption levels. The plots clearly indicate that until 17% corruption, the validation accuracy does not reduce significantly. But, beyond 17% corruption level, the model begins to show instability in the validation accuracies. Further, we observe that the model is able to avoid overfitting, as with the increase in the corruption level, the training accuracy falls, but never the validation accuracy.
@@
Figure 10: Qualitative results showing the effect of corruption level on the ground truth (second column from the left) and the final predicted (rightmost column) mask, for the input image shown in Figure 3(a) (leftmost column). The predicted masks prior to applying threshold to the U-Net model outcomes (second column from the right) get blurrier with increase in corruption, but are robust in capturing river structure in the presence of morphological label noise.
@@
Figure 11: Qualitative results showing the effect of corruption level on the ground truth (second column from the left) and the final predicted (rightmost column) mask, for the input image shown in Figure 3(b) (leftmost column). The predicted masks prior to applying threshold (second column from the right) to the U-Net model outcomes get blurrier with increase in corruption, but are robust in capturing river structure in the presence of morphological label noise.
@@
Figure 12: Results showing the effect of corruption level on the ground truth (second column from the left) and the final predicted (rightmost column) mask, for the input (leftmost column). The predicted masks prior to applying threshold (second column from the right) to the U-Net model outcomes capture the river geometry, but the final outputs indicate the model’s failure.
@@
These results are further supported by Figures 10 and 11. These figures show the effect of corruption of training data on two examples. They show the input image, and the corresponding actual masks which are fed to the model as inputs. These are points on which erosion has been applied, and we observe that with the increase in the corruption levels, the percentage of white pixels reduces significantly, with no remaining white pixels when the corruption is increased to 25%. This is also an example of a data point where just increasing the corruption level does not necessarily lead to an increase in the actual corrupted pixels.
@@
With the increase in corruption in the training data, we observe that the baseline U-Net model output, that is gotten before applying the threshold of 0.5 (explained in Section 3.2), becomes increasingly blurred. This indicates that the increase in corruption reduces the model’s ability to discern land from water. However, the model still remains capable of finding the river structure or geometry, and when we apply the threshold on the U-Net output, the masks predicted are still reasonably accurate. This indicates that U-Net is robust to high levels of corruption. These observations are shown in Figures 10 and 11, where the model predictions at 20% and 25% corruption cases are correct in terms of the river structure, but the overall accuracy has reduced in the U-Net model outputs, obtained before the step of applying the threshold of 0.5.
@@
We also include the results from a data point on which the clean U-Net model itself does not perform very accurately. One such failure case is shown in Figure 12. The model faces issues in accurately segmenting narrow tributaries in the image, and the model performance further deteriorates with the increase in corruption. This example also appears to show that corruption with dilation is much more severe than erosion. This observation is also supported by Figure 8, which shows that for dilation, the number of corrupted pixels continues to increase with the increase in corruption level, unlike the case with erosion. The prediction is poor even at 10% corruption.
@@
Figure 14: A data point where the ground truth is incorrect due to incomplete ground truth owing to unavailability of Sentinel-2 image in the specific region.

### v2

@@ 1 Introduction
Water body segmentation is a critical task in remote sensing and geospatial analysis, which caters to diverse and critical applications such as flood mapping and monitoring [1], glacial lake segmentation [2], sedimentation and erosion studies [3], and environmental impact assessments. These tasks often rely on high-resolution satellite imagery, such as Synthetic Aperture Radar (SAR), and consequently, accurate segmentation models to delineate water bodies from surrounding land features. SAR imaging uses energy reflectance to be available through day-night lighting and all-weather conditions. However, SAR sensors do not capture data that can discriminate between water and water-like surfaces [4]. Inland water body segmentation is far more challenging than coastline detection, when discussing land-water segmentation of SAR images, owing to the complex geometry of hydrological structures, such as river basins with several tributaries and distributaries (Figure 1).
@@
Deep learning (DL) models, namely, convolutional neural networks (CNNs), outperform traditional approaches in SAR image segmentation in several scenarios. Particularly, encoder-decoder architectures in CNNs, such as U-Net [5], have demonstrated state-of-the-art performance in land-water segmentation of SAR images. Despite these advances in DL, water body segmentation models continue to face challenges during the learning process, such as the presence of noise, domain shifts in remote sensing data, and the impact of adversarial perturbations. Several works in this area rely primarily on manual annotation to generate ground truth masks [1, 6, 7]. These are prone to several human errors, especially during manual annotation. These inadvertent errors include incorrect labels due to variations in river width and mislabeling boundary pixels. These errors tend to propagate to the image segmentation models using DL, which calls for testing the robustness of these architectures. This motivates our investigation into potential labeling errors and their simulation. Our approach in this study is the novel usage of morphological noise, followed by a systematic testing of adversarial learning of segmentation models, specifically the U-Net architecture. Morphological noise closely models the incorrectness in labels owing to the complex geometry of the hydrological structures of a large river basin, e.g., the Amazon River Basin in South America.
Adversarial attacks pose a significant threat to the robustness and reliability of machine learning models. These attacks involve intentionally crafted perturbations to input data that cause models to misclassify or produce incorrect outputs, often with changes that are barely discernible to the human eye. In the context of water body segmentation, adversarial attacks can lead to inaccurate delineation of water boundaries, with potentially severe consequences for downstream applications such as flood risk assessment or glacial melt monitoring [8, 9]. In this work, the primary form of corruption that we experiment with is data poisoning attacks, which involve corruption of the training data passed to the model. We introduce a specific type of noise, which is the morphological label noise, that is generated using morphological operations (erosion and dilation) in a controlled manner to introduce corruptions in the data points, thus simulating boundary errors.
@@ 2 Related Work
The identification of water bodies from Earth Observation (EO) images is an important step in many applications. Conventionally, a thresholding-based approach is used for water bodies identification, but the thresholds are identified experimentally based on a subset of data [13]. This lacks the capability to generalize, and thresholds may be different for different subsets. Tools based on spectral indices for river identification [14], along with others based on a mathematical morphological-based approach, have also been proposed [15]. However, determining the parameters of path opening and closing is challenging and can affect the results obtained.
@@
Recently, semantic segmentation-based approaches have been utilized to identify water bodies in EO images. Semantic segmentation aims to assign a class label to each pixel in an image. In the past decade, deep learning-based approaches have obtained excellent results in semantic segmentation for various applications. For instance, several deep learning based approaches have been utilized for semantic segmentation of water bodies [1] [16], including the usage of networks such as FCN-8s [17], tiramisu [18], and pix2pix [19], which is a GAN which has been adapted for the segmentation task. In these supervised learning based segmentation approaches, the training images must be manually labelled by a human. This is a time-consuming and error-prone task. Unsupervised learning methods for segmentation on Sentinel-1 SAR imagery [20] have also been explored for river identification. Other interesting works include a Dynamic Snake Convolution [21], which can be used to segment tubular structures such as blood vessels and roads and may be adapted for river segmentation. However, these segmentation approaches focus mainly on accurately identifying water bodies. In comparison, the proposed work studies the performance of segmentation models (primarily U-Net) in water body identification under adversarial attacks from EO images. Our work for semantic segmentation is inspired by DeepRivWidth [22], which uses two primary segmentation models, namely U-Net [5] and DeepLabV3+ [23] for segmentation of Sentinel-1 SAR images.
@@
There exist multiple datasets for water body identification, along with various sources in the form of shapefiles for water body mapping [24, 25, 26, 27, 28]. Examples include the Sen1Floods11 dataset [24], which uses raw Sentinel-1 imagery, and classifies flood water and permanent water, for flood mapping at a global scale, and the Hydrosheds Dataset [25], which provides hydrographic data products including catchment boundaries, rivers, lakes, etc. Other examples include DIVA-GIS [26], which provides comprehensive shapefiles for river basins that can be used for ground truth generation, and MERIT Hydro [27], which provides high-resolution hydrography maps using elevation data and water body datasets such as G1WBM (Global 1-second Water Body Map), GSWO (Global Surface Water Occurrence), and OpenStreetMap.
@@
Many of the previously discussed works use human annotation to generate the ground truth for all the images. Not only is this a laborious and time-consuming task, but it is also highly error-prone. There have been several works aimed at studying the error types and their effects on the model performance. Most errors from a human annotator occur at the physical boundaries of class labels in spatial applications, which may be classified as biased or unbiased errors [29]. Biased errors may be repeated under similar conditions by the annotator due to human limitations in comprehension and discernibility of the classes, while unbiased errors stem from circumstances (such as a slip of a hand), and are less likely to be repeated in relabeling.
@@
Adversarial attacks are generally classified into white- and black-box attacks, based on the level of access granted to the attacker [32]. In white-box attacks, the attackers obtain all information about the target model, including its structure, parameters, etc, and they generally directly interact with the model. On the other hand, for black-box adversaries, the attacker only needs access to the output information of the target model to launch an attack. Black-box attacks are generally more realistic attacks, as the adversary may not always have access to the internal workings of the system. Data poisoning attacks can be either black- or white-box, as they involve changes to the training data itself. Also, depending on the nature of the changes made, they can be considered either a white- or a black-box change [33].
@@
There have also been several works that specifically target data poisoning attacks. One such work proposes a poisoning technique similar to ours, where the adversary targets labels in a multi-class setting and is constrained under a label-flipping budget [36], which is an idea we employ in our work as well. There also exist targeted poisoning attacks, which target one specific test instance, rather than the entire training set [37]. Our work targets the training data instances via a controlled corruption mechanism, which uses morphological operations, and hence, is different from previous studies.
@@ 3 Method
Figure 3: Examples of Sentinel-1 SAR images of the Amazon River Basin (left) in our dataset, along with their generated binary water masks (right).
@@ 3.1 Dataset Generation
We selected Sentinel-1 SAR data for our case study of the Amazon River Basin from platforms like the Sentinel Scientific Data Hub and Copernicus Open Access Hub. The Amazon River Basin is an area of interest due to the large area coverage of 7,000,000 km2, with a widespread complex river system. These river structures provide sufficient variety in the training and testing data, as shown in Figure 1. Also, the physical structures of the river and its channels are of complex geometry, which provides several regions where human error could occur while annotating. The chosen images cover 25,194.73 km2 of the basin. The geographic locations of the images in our dataset are shown in Figures 4 and 5.
@@
    The first step is the Orbit File Application.
@@
Finally, we implemented a widely-adopted approach using the NDWI index [28]. The multispectral images from Sentinel-2 were obtained using Google Earth Engine. For each image, we clip out the region corresponding to the range of latitude and longitude covered by each image, and then, we calculate the NDWI index. This index is used to identify the water regions from the land regions, using a suitable threshold, for each image. This resulted in highly accurate water masks. Examples of the SAR images and their corresponding masks are shown in Figure 3. The entire curated dataset (with correct ground truths and adversarial examples) is published at https://figshare.com/articles/dataset/Adversarial_Attack_Erosion_and_Dilation_Dataset_for_Amazon_River_Basin/28784405?file=53625587.
@@ 3.2 Models for Segmentation
We trained a baseline model using a vanilla U-Net [5] architecture, using our dataset with the accurate, i.e., uncorrupted, ground truth labels. Our goal with this initial setup was to establish a baseline model to benchmark the performance of the model on our dataset. The U-Net model has an encoder with four levels, starting with 64 filters and doubling the number at each level, reaching 512 filters. After downsampling, the bottleneck layer has two convolutional layers with 1024 filters and a dropout layer. The decoder then upsamples the feature maps back to the original resolution. The decoder also applies residual convolutional blocks to refine feature extraction. We then apply a simple threshold of 0.5 on the output of U-Net to segregate land and water pixels and generate a binary mask.
@@ 3.3 Simulation of Adversarial Attacks in Training
We also introduce randomness in the kernel sizes used for corruption, to simulate the real-world non-uniformity in the corruption. The stochastic process of selecting the kernels is essential when we need the model not to be biased in its learning. The chosen size depends on the percentage of white pixels (the pixels with mask value 1) in the image. For the images that have a white pixel percentage below 0.2, we give higher probability for choosing the kernel of size 7$\times$7 and lower probability for those with 5$\times$5 and 3$\times$3. We give the probability of 0.7 to that kernel and 0.15 to each other kernel. The exact opposite happens for dilation, we give high probability for 3$\times$3 and lower probability values for the other two kernels when the white pixel percentage is high. Finally, we utilize the following levels of corruption, given the image sizes are $512\times 512$: $2\%$, $12\%$, $15\%$, $17\%$, $20\%$, $25\%$, $30\%$}.
With the model trained on correct masks, we now introduce controlled corruptions to the masks by using erosion and dilation, as these operations present examples of potential human errors in the precise delineation of land-water boundaries. Thus, these controlled corruptions are morphological label noise, and this process of corruption is effectively a data poisoning procedure.
@@ 3.4 Implementation
The data points are divided into 883 training, 127 validation, and 253 test samples. The U-Net model is trained for 20 epochs, for this experiment and across all subsequent experiments. We used a pre-trained implementation of DeepLabV3 with ResNet-50 backbone, and finetuned it over our dataset, for 50 epochs (with an Early Stopping condition to prevent overfitting). We used SegFormer (b-4-sized) pretrained on Imagenet [46], and finetuned for 50 epochs on our dataset. Finally, for Mask2Former, we used the large-sized version with Swin Transformer, pretrained on the Cityscapes dataset [47], and finetuned it for a total of 50 epochs on the dataset. We run all the models on a system with an AMD Ryzen 9 5950X 16-core processor, connected to an NVIDIA RTX A4000 graphics card. Our code is available at https://github.com/GVCL/IWSeg-SAR-Poison.git.
@@ 4.1 Segmentation Results with the Uncorrupted Dataset
We present the results of state-of-the-art segmentation models (U-Net, DeepLabV3, SegFormer, Mask2Former) on our dataset (uncorrupted), which helps us assess the quality of the dataset. Table 1 shows the performance of the models on the dataset. The performance of each of the models saturates at a per-pixel accuracy of $\sim$0.955. Moreover, the IoU for all these models is comparable.
We further investigate the samples that none of the models are able to segment properly. A few such failure cases are discussed in Section 4.3, where either due to incorrect thresholding, or due to unavailability of a complete Sentinel-2 image, the ground truth itself is incorrect. The incorrectness of the ground truth consequently leads to a higher level of segmentation inaccuracy.
However, one consistent failure observed across the model is its inability to fully delineate narrow tributaries in the image. An example of one such failure case is shown in Figure 6. Overall, we observe that all our selected models face issues in delineating the very narrow tributaries and are only somewhat correct, with varying levels of success.
@@ 4.2 Adversarial Experimentation
This saturation of corruption is demonstrated in Figures 8 and 8. Figure 8 shows the box plot of the actual number of corrupted pixels across all training images for different corruption levels, where the corruption is by erosion, and a similar box plot for dilation. We use the Structural Similarity Index (SSIM) score to quantify the similarities between two images. Figure 8 shows the distribution of the SSIM scores between the uncorrupted ground truth and the corresponding corrupted ground truth, computed across all images and for different corruption levels. We observe that the number of pixels corrupted for images undergoing erosion becomes stagnant as we increase the corruption levels, with the median and the lower quartile barely changing beyond $20\%$ corruption. The dilation plots, on the other hand, show the expected trends. The same is true for the SSIM values between the uncorrupted masks and their corresponding corrupted masks. Thus, we conclude that, in our approach, the corruptions may be limited by design, as the final count of corrupt pixels for an image may be less than that expected in the targeted level of corruption.
@@ 4.3 Results on Corrupted Dataset and Discussion
Table 2 shows the evaluation metrics of training the U-Net model across the various levels of corruption. We get a good 95% accuracy in segmentation. Furthermore, the performance is mostly unaffected even after increasing the corruption levels to as high as 20%. This is surprising, as the model is able to tolerate even such high levels of corruption. We observe that the model performance does not deteriorate significantly until 25% corruption. Hence, we investigate more into the training of the model, the qualitative predictions of the model on a particular image, and the actual corruption that the model is facing.
@@
Figure 9 shows the per-epoch accuracy on the training and validation data for each of the corruption levels. The plots clearly indicate that until 17% corruption, the validation accuracy does not reduce significantly. But, beyond a 17% corruption level, the model begins to show instability in the validation accuracies. Further, we observe that the model is able to avoid overfitting, as with the increase in the corruption level, the training accuracy falls, but never the validation accuracy.
@@
Figure 10: Qualitative results showing the effect of corruption level on the ground truth (second column from the left) and the final predicted (rightmost column) mask, for the input image shown in Figure 3(a) (leftmost column). The predicted masks before applying the threshold to the U-Net model outcomes (second column from the right) get blurrier with an increase in corruption, but are robust in capturing river structure in the presence of morphological label noise.
@@
Figure 11: Qualitative results showing the effect of corruption level on the ground truth (second column from the left) and the final predicted (rightmost column) mask, for the input image shown in Figure 3(b) (leftmost column). The predicted masks before applying the threshold (second column from the right) to the U-Net model outcomes get blurrier with an increase in corruption, but are robust in capturing river structure in the presence of morphological label noise.
@@
Figure 12: Results showing the effect of corruption level on the ground truth (second column from the left) and the final predicted (rightmost column) mask, for the input (leftmost column). The predicted masks before applying the threshold (second column from the right) to the U-Net model outcomes capture the river geometry, but the final outputs indicate the model’s failure.
@@
These results are further supported by Figures 10 and 11. These figures show the effect of corruption on training data in two examples. They show the input image and the corresponding actual masks, which are fed to the model as inputs. These are points on which erosion has been applied, and we observe that with the increase in the corruption levels, the percentage of white pixels reduces significantly, with no remaining white pixels when the corruption is increased to 25%. This is also an example of a data point where just increasing the corruption level does not necessarily lead to an increase in the actual corrupted pixels.
@@
With the increase in corruption in the training data, we observe that the baseline U-Net model output, which is obtained before applying the threshold of 0.5 (explained in Section 3.2), becomes increasingly blurred. This indicates that the increase in corruption reduces the model’s ability to discern land from water. However, the model remains capable of finding the river structure or geometry, and when we apply the threshold on the U-Net output, the masks predicted are still reasonably accurate. This indicates that U-Net is robust to high levels of corruption. These observations are shown in Figures 10 and 11, where the model predictions at 20% and 25% corruption cases are correct in terms of the river structure, but the overall accuracy has reduced in the U-Net model outputs, obtained before the step of applying the threshold of 0.5.
@@
We also include the results from a data point on which the clean U-Net model itself does not perform very accurately. One such failure case is shown in Figure 12. The model faces issues in accurately segmenting narrow tributaries in the image, and the model’s performance further deteriorates with the increase in corruption. This example also appears to show that corruption with dilation is much more severe than erosion. This observation is also supported by Figure 8, which shows that for dilation, the number of corrupted pixels continues to increase with the increase in corruption level, unlike the case with erosion. The prediction is poor even at 10% corruption.
@@
Figure 14: A data point where the ground truth is incorrect due to incomplete ground truth owing to the unavailability of the Sentinel-2 image in the specific region.